MICROWAVE STUDIES OF PERFLUOROPENTANE AND ITS HELICITY

\underline{JOSEPH A. FOURNIER}, ROBERT K. BOHN, Dept. of Chemistry, Univ. of Connecticut, Storrs, CT 06269-3060; JOHN A. MONTGOMERY, JR., Dept. of Physics, Univ. of Connecticut, Storrs, CT 06269-3046.

Unlike hydrocarbons, fluorocarbons are helical in their all-trans most stable conformations. A definitive experimental determination of the helical angle has not been accomplished because single crystal X-ray studies are not available due to the lack of suitable crystallization solvents, but a value of 17^{∘} from exactly trans is generally accepted from X-ray fiber studies. Using a pulsed-jet Fourier Transform microwave spectrometer, we have observed and assigned the rotational spectra of the lowest energy all-anti C_{2} symmetry form of perfluoro-n-pentane and all three of its ^{13}C isotopomers. A, B, and C values of the parent species are 990.6394(4) MHz, 314.00020(14) MHz, and 304.37034(14) MHz, respectively. A range of effective r_{0} structures incorporating various model constraints are consistent with about 16^{∘} torsion and a Kraitchman analysis of the parent and ^{13}C species gives 13.3^{∘}. Ab initio calculations are consistent with the experimental results.